# Can Large Language Models Understand Symbolic Graphics Programs?

### v1

@@ Appendix B Human Study Details
We ran a human study to verify the labels produced by GPT-4o for the benchmark over a subset of $500$ stimuli. We recruited $55$ participants from the crowdsourcing platform, Prolific [71]. Stimuli were batched into $10$ sets of $50$ stimuli each. Each participant was randomly assigned a batch of $50$ stimuli; stimuli were presented in a random shuffled. On each trial, participants saw the question, original image, and set of multiple choice options. Participants selected an option by clicking a button. We include an example screenshot of a trial in Figure 11. Participants were paid at a base rate of \$$\$12.50$/hr. They were informed that they could receive a bonus up to \$$\$15/hr$ based on the amount of correct answers they achieved. All participants received the full bonus. Our study was approved by our institutional ethics review board, and all participants provided informed consent. We include the set of instructions and sample screenshots in Figures 12 and 13, respectively. We found high inter-annotator agreement (participants in the same batch had between $0.7-0.85$ Fleiss Kappa’s alpha agreement, where higher implies higher agreement). We find that humans’ mode response matched GPT-4o on 90% of the examples (450 of the 500 stimuli).
@@ D.2 CAD Data
Figure 21: CAD examples (3Drecon.) in our SGP-Bench.
@@ F.1 Template for benchmark construction
"Describe in detail the semantic or geometric characteristics
of the object shown in the image."
@@
"Can you provide a detailed account of the geometric or
semantic features of the object in the image?"
@@
"Could you detail the geometric or semantic features of the
object in the image?"
@@
"Please describe the semantic or geometric features of the
object in the image comprehensively."
@@
"Can you elaborate on the semantic or geometric features of
the object in the image?"
@@
"Explain the semantic or geometric features of the object in
the image in detail."
@@
"Elaborate on the semantic and geometric characteristics of
the object shown in the image."

### v3

@@ Appendix B Human Study Details
We ran a human study to verify the labels produced by GPT-4o for the benchmark over a subset of $500$ stimuli. We recruited $55$ participants from the crowdsourcing platform, Prolific [73]. Stimuli were batched into $10$ sets of $50$ stimuli each. Each participant was randomly assigned a batch of $50$ stimuli; stimuli were presented in a random shuffled. On each trial, participants saw the question, original image, and set of multiple choice options. Participants selected an option by clicking a button. We include an example screenshot of a trial in Figure 14. Participants were paid at a base rate of \$$\$12.50$/hr. They were informed that they could receive a bonus up to \$$\$15/hr$ based on the amount of correct answers they achieved. All participants received the full bonus. Our study was approved by our institutional ethics review board, and all participants provided informed consent. We include the set of instructions and sample screenshots in Figures 15 and 16, respectively. We found high inter-annotator agreement (participants in the same batch had between $0.7-0.85$ Fleiss Kappa’s alpha agreement, where higher implies higher agreement). We find that humans’ mode response matched GPT-4o on 90% of the examples (450 of the 500 stimuli).
@@ D.2 CAD Data
Figure 24: CAD examples (3Dcomplex) in our SGP-Bench.
@@ F.1 Template for benchmark construction
"Describe in detail the semantic or geometric characteristics of
the object shown in the image."
@@
"Can you provide a detailed account of the geometric or semantic
features of the object in the image?"
@@
"Could you detail the geometric or semantic features of the object
in the image?"
@@
"Please describe the semantic or geometric features of the object
in the image comprehensively."
@@
"Can you elaborate on the semantic or geometric features of the
object in the image?"
@@
"Explain the semantic or geometric features of the object in the
image in detail."
@@
"Elaborate on the semantic and geometric characteristics of the
object shown in the image."